# Exploring Partial Multi-Label Learning via Integrating Semantic Co-occurrence Knowledge

### v1

@@ Abstract
Partial multi-label learning aims to extract knowledge from incompletely annotated data, which includes known correct labels, known incorrect labels, and unknown labels. The core challenge lies in accurately identifying the ambiguous relationships between labels and instances. In this paper, we emphasize that matching co-occurrence patterns between labels and instances is key to addressing this challenge. To this end, we propose Semantic Co-occurrence Insight Network (SCINet), a novel and effective framework for partial multi-label learning. Specifically, SCINet introduces a bi-dominant prompter module, which leverages an off-the-shelf multimodal model to capture text-image correlations and enhance semantic alignment. To reinforce instance-label interdependencies, we develop a cross-modality fusion module that jointly models inter-label correlations, inter-instance relationships, and co-occurrence patterns across instance-label assignments. Moreover, we propose an intrinsic semantic augmentation strategy that enhances the model’s understanding of intrinsic data semantics by applying diverse image transformations, thereby fostering a synergistic relationship between label confidence and sample difficulty. Extensive experiments on four widely-used benchmark datasets demonstrate that SCINet surpasses state-of-the-art methods.
Index Terms:
Semantic co-occurrence, multi-modal representation, partial multi-label learning.
@@ I Introduction
Multi-label learning has demonstrated tremendous potential in fields. However, due to the high cost of labeling and the subjectivity of annotators, real-world datasets often suffer from incomplete and noisy labels. This challenge has spurred the exploration of partial multi-label learning methods aimed at addressing these issues more effectively. Consequently, driven by this research need, partial multi-label learning has garnered vibrant attention in machine learning [1, 2]. It represents a new paradigm for multi-label recognition (MLR) and has been widely applied in micro-video classification [3], image recognition [4]. In these applications, the training dataset typically contains only partial labels rather than complete ones, which better reflects the complexity of real-world data. For instance, in a scenario where one instance corresponds to a single label, the model only needs to identify the correct label for that specific instance (Fig. 1 (a)). Conversely, in a setting where multiple instances correspond to a single label, the model must extract features from several instances to determine the common label, (Fig. 1 (b)).
@@
The key to addressing partial multi-label learning tasks lies in deeply mining discriminative semantic features and constructing a collaborative representation model to fully leverage both label and instance information. A straightforward approach involves leveraging matrix factorization with auxiliary information to accelerate label completion and enhance prediction performance[5]. However, such works essentially suffer from performance limitations due to their neglect of high-order correlations among labels. This inherent constraint has motivated extensive research on capturing label correlations and exploiting data structural information through various mechanisms. For instance, Xie et al. [6] proposed a partial label framework that assigns confidence scores to candidate labels while utilizing data structural information to jointly optimize the classification model and refine confidence estimations for ground-truth labels. Concurrently, semantic-aware modules [7, 8] and loss calibration techniques [9] have been widely adopted in partial multi-label frameworks due to their competitive performance in exploring label correlations.
@@
To date, visual-language pre-trained models have exhibited outstanding performance across various visual tasks. These models, such as the CLIP [12], pre-trained on a vast corpus of image-text pairs, benefit from large-scale pre-training and effectively bridge the gap between vision and language, providing rich prior knowledge for downstream tasks. In the context of partial multi-label learning, these pre-trained models have demonstrated significant advantages. Even when annotations are available for only a subset of categories, they can leverage their generalization capabilities and existing knowledge of visual-language associations to accurately infer unseen categories, thereby enhancing recognition accuracy and robustness. Sun et al. [13] utilized the strong alignment capability of pre-trained visual-language models by introducing a pair of learnable positive and negative prompts, along with a category-specific region feature aggregation method, addressing how to train models and improve recognition accuracy under limited annotations effectively. Ding et al. [14] particularly highlighted the importance of leveraging structured semantic relationships between labels to address the lack of supervision in situations of incomplete labeling. By introducing cross-modal prompters and a semantic association module, their method more effectively extracts and utilize the rich knowledge embedded in pre-trained visual-language models, such as CLIP, resulting in performance improvements on MLR tasks beyond existing approaches. In addition, Wang et al. [15] proposed a hierarchical semantic prompting network using CLIP for hierarchical prompt learning, aimed at multi-label classification tasks with a single positive label. As a result, we have adopted techniques based on multi-modal large language models to tackle partial multi-label learning tasks, thereby enhancing performance and accuracy.
@@
Although the aforementioned methods have achieved satisfactory results to a certain extent, their effectiveness remains constrained in addressing the challenge of insufficient supervision, particularly regarding inadequate exploitation of supervisory signals among instances. In reality, inter-instance relationships can provide crucial clues for understanding and predicting attributes of unlabeled instances. Taking image data as an example, instances within images typically do not exist in isolation but exhibit intrinsic interconnections. Such relational patterns offer additional contextual information that facilitates more accurate target recognition under incomplete label supervision. Moreover, large-scale cross-modal datasets inherently contain rich relational information among instances. While most current multi-modal representation learning approaches have yet to fully exploit these patterns and interactions in partial multi-label learning tasks, they demonstrate significant potential for enhancing model capabilities in comprehending and modeling instance-level associations within complex scenes.
@@
Based on the aforementioned issues, we propose a novel approach to address these challenges. This method begins by deeply excavating the discriminative semantic features embedded within the image context, and subsequently constructs a collaborative representation model of labels and instances in a high-dimensional data space, thereby achieving precise mapping across a cross-modal semantic space. Specifically, to tackle the challenges posed by incomplete labels, we focus on matching of multiple semantic co-occurrence relationships between labels and instances. By thoroughly analyzing their interaction patterns, the model is capable of not only identifying existing semantic labels within the data but also inferring unseen labels based on the associations among known instances. Consequently, this significantly enhances the overall performance of partial multi-label learning tasks in complex scenarios and large-scale cross-modal environments, offering a new perspective on addressing the challenge of partial labeling.
@@ II Related Work
This section provides a concise overview of previous research on multi-label image classification tasks, focusing on MLR with comprehensive annotations, partial multi-label learning, and multi-modal representation learning in visual tasks.
@@ II-B Partial Multi-Label Learning
In partial multi-label learning, each instance is associated with only a subset of its true labels, with the remaining unobserved. This scenario is more challenging than traditional multi-label learning with complete label assignments. Formally, given a training dataset $\mathcal{D}=\{(\mathbf{X}_{i},Y_{i})\}_{i=1}^{N}$, where $\mathbf{X}_{i}\in\mathbb{R}^{d}$ is a d-dimensional feature vector representing the $i$-th image, and $Y_{i}\subseteq\mathcal{Y}=\{\mathbf{l}_{1},\mathbf{l}_{2},\cdots,\mathbf{l}_{q}\}$ is the set of labels associated with the $i$-th image, with $\mathcal{Y}$ being the complete set of possible labels containing $q$ labels. Early studies [23] often treated missing labels as negative labels by default. However, this simplistic approach can introduce significant label bias, as it may erroneously classify a substantial number of true positive labels as negative, thereby adversely affecting the model’s learning performance. To address this issue, several methods have been proposed. For instance, some approaches leverage instance similarity, label co-occurrence information, and low-rank regularization to recover missing labels through matrix completion techniques [24]. Kapoor et al. [25] treated missing labels as latent variables within a probabilistic framework and updated them via Bayesian inference. Additionally, Wu et al. [26] introduced a third state for missing labels and reconstructed the complete label information by enforcing consistency constraints and label assignment smoothness constraints. It is noteworthy that some earlier works [27] still adhered to the strategy of treating unknown labels as negative, while the approach of decomposing the multi-label task into multiple independent binary classification problems overlooked the inherent dependencies among labels.
@@
In light of the aforementioned limitations, subsequent research has increasingly focused on leveraging label correlations to supplement unknown label information, employing methods such as low-rank regularization, label dependency propagation, and hybrid graph modeling [5]. However, these approaches typically require loading the entire training dataset into memory, making them difficult to integrate with fine-tuning strategies for deep neural networks based on mini-batch training. To address this, recent studies have shifted their attention to designing novel loss functions and optimization strategies. For instance, Durand et al. [4] proposed a generalization of the standard binary cross-entropy loss by incorporating label proportion information. Huynh and Elhamifar [28] introduced regularization to the loss function by considering statistical co-occurrence and image-level feature similarity. Pu et al. [8] enhanced samples by generating composite labels through the mixing of category-specific features in the feature space. Meanwhile, Xie et al. [29] utilized structural information from both feature and label spaces to assign confidence scores to potential labels, while Wang et al. [30] guided label disambiguation using interior-point methods and gradient boosting algorithms. He et al. [31] designed a soft labeling threshold operator to narrow the gap between candidate labels and ground truth labels. On the other hand, Guan et al. [32] proposed a semi-supervised partial multi-label learning method that combines partially labeled data with unlabeled data, capturing intrinsic label relationships through low-rank and manifold constraints. To address the issue of significant dependencies among labels, Chen et al. [7] inferred missing labels by leveraging known label information within and across images, leading to the development of the HST method [16], which is based on category-specific feature-prototype similarity and differential threshold learning. Although HST incorporates adaptive threshold learning, it still relies on modeling inter-class similarities, which may vary across datasets, potentially affecting generalization performance. These methods primarily rely on co-occurrence information within datasets, leading to suboptimal inference when labels are sparse or unstable. In contrast to existing approaches, our method leverages pre-trained multi-modal models, enabling effective reasoning in low-label scenarios by incorporating prior knowledge. To tackle the issue of incomplete label annotations, we introduce an intrinsic semantic augmentation strategy to ensure robust performance.
@@ II-C Multi-modal Semantic Fusion
Based on this context, researchers have proposed various innovative methods to further explore and leverage the semantic relationships between vision and language. For instance, Wang et al. [15] developed a novel hierarchical semantic prompting network that dynamically captures the hierarchical semantic relationships between images and text, effectively enhancing the model’s ability to understand complex semantic structures. Concurrently, He et al. [36] introduced an open-vocabulary multi-label classification framework based on multi-modal knowledge transfer, aiming to address the issue of unseen labels in real-world MLR systems. Additionally, Wang et al. [37] designed a multi-modal context prompting learning method, which incorporates learnable context prompts in both image and text modalities and utilizes coupling functions to facilitate interactive connections between modal prompts, thereby strengthening the correspondence between images and labels and improving the accuracy of MLR. Liu et al. [38] proposed a novel multi-label few-shot image classification framework based on paired feature enhancement and flexible prompt learning to explore further and leverage the semantic relationship between vision and language. Rawlekar et al. [39] introduced PositiveCoOp, which learns only positive prompts and replaces negative prompt learning with embedded vectors learned directly in the shared feature space without relying on a text encoder. Liu et al. [40] presented a language-driven cross-modal classifier, which trains a classifier using language data generated by large language models and then transfers it to the visual modality. Fang et al. [41] proposed ProPOT, which generates improved class prototypes by partially aggregating local image features through a partially optimal transport scheme.
Despite these advancements, existing multi-modal learning methods still exhibit limitations. They fail to adequately capture the complex co-occurrence relationships between labels and instances, often overlooking cross-dependencies, which adversely affects alignment performance. Moreover, current models tend to focus on local similarities while lacking the integration of global correlations, thereby constraining the accurate estimation of label confidence. To address these issues, our approach employs global modeling of labels, instances, and their multiple co-occurrence relationships, achieving multi-level semantic fusion and significantly enhancing alignment precision. Furthermore, we construct a cross-modal fusion module to deeply integrate textual and visual data, optimizing label confidence estimation. To tackle the problem of incomplete label annotations, we introduce an intrinsic semantic augmentation strategy, ensuring robust performance.
@@ III Methodology
We propose to learn and infer complex semantic co-occurrence relationships between the visual and textual feature spaces, particularly among labels and labeled instances. This approach enhances the model’s proficiency in identifying rare or ambiguous categories. It boosts its generalization skills, enabling it to navigate the complexity and diversity of the real world more effectively. Fig. 2 outlines our proposed SCINet method. Initially, we process and interact with input data using the bi-dominant prompter module. Simultaneously, Learnable Prompts acquire textual labels in the form of a vector sequence, providing context for the given label names. Subsequently, text-dominant and image-dominant encoders process textual and image features, respectively. The cross-modality fusion module then deeply combine text and image information to understand and classify data more effectively. Furthermore, we utilize an intrinsic semantic augmentation strategy to enable the model to comprehend data features and encourage a synergistic relationship between label confidence and sample complexity.
@@ III-A Bi-Dominant Prompter
Owing to the advancements in large-scale pre-trained models, such as BERT and CLIP, we now have the capability to effortlessly access contextual embeddings of labels. This advancement enables the effortless inference of co-occurrence relationships or correspondences between labels and instances. Such strategy is especially attractive when adequate label supervision is lacking. Furthermore, the extensive prior knowledge contained within pre-trained models facilitates the connection between known and unknown labels, potentially alleviating the challenges associated with limited label supervision. Therefore, we introduce a semantics-driven prompter to explore the label-to-label co-occurrence relationships inherent in the pre-trained model. Given the widespread attention and outstanding performance of vision-language models in computer vision, we have chosen CLIP as the primary focal of our research. Let $V=[v_{1},v_{2},\cdots,v_{m},\text{CLS}]$, where each $v_{i}$ represents a learnable soft prompt token, and CLS denotes the given label name. The entire set of labels is denoted as $Y$, representing each instance as $y_{i}\in Y$. The text-dominant encoder accepts label information while simultaneously processing the image set. The textual representation obtained from the text-dominant encoder process is denoted as $z$. The input image $\mathbf{X}$ is represented as $\mathbf{X}=\left\{\left(\mathbf{s_{1}},C_{1},U_{1}\right),\left(\mathbf{s_{2}},C_{2},U_{2}\right),\cdots,\left(\mathbf{s_{n}},C_{n},U_{n}\right)\right\}$, where $s_{i}$ denotes the $i$-th instance, $C_{i}=\{\mathbf{c}_{i1},\mathbf{c}_{i2},\cdots\}$ represents its set of known correct labels, and $U_{i}=\{\mathbf{u}_{i1},\mathbf{u}_{i2},\cdots\}$ corresponds to its set of unknown labels. To enhance the model’s robustness, three different levels of transformations are applied to the input, including weak transformation ($X^{-}$), original transformation ($X$), and strong transformation ($X^{+}$). These three transformations are collectively integrated as $X^{{}^{\prime}}$ and utilized for subsequent learning and optimization processes. For a detailed explanation of the three different levels of transformation, please refer to Section III-C. The visual representation obtained, denoted as $f$, is extracted through the modified CLIP image encoder. This dual-encoder approach enables the model to understand and apply the semantic co-occurrence knowledge, thereby improving the performance of MLR. Specifically, the text-dominant encoder and the visual-dominant encoder are defined as follows:

### v2

@@ Abstract
Partial multi-label learning aims to extract knowledge from incompletely annotated data, which includes known correct labels, known incorrect labels, and unknown labels. The core challenge lies in accurately identifying the ambiguous relationships between labels and instances. In this paper, we emphasize that matching co-occurrence patterns between labels and instances is key to addressing this challenge. To this end, we propose the Semantic Co-occurrence Insight Network (SCINet), a novel and practical framework for partial multi-label learning. Specifically, SCINet introduces a bi-dominant prompter module, which leverages an off-the-shelf multimodal model to capture text-image correlations and enhance semantic alignment. To reinforce instance-label interdependencies, we develop a cross-modality fusion module that jointly models inter-label correlations, inter-instance relationships, and co-occurrence patterns across instance-label assignments. Moreover, we propose an intrinsic semantic augmentation strategy that enhances the model’s understanding of intrinsic data semantics by applying diverse image transformations, thereby fostering a synergistic relationship between label confidence and sample difficulty. Extensive experiments on four widely-used benchmark datasets demonstrate that SCINet surpasses state-of-the-art methods.
@@ I Introduction
Multi-label learning has demonstrated tremendous potential in various fields. However, due to the high cost of labeling and the subjectivity of annotators, real-world datasets often suffer from incomplete and noisy labels. This challenge has spurred the exploration of partial multi-label learning (PML) methods, which aim to address these issues more effectively. Consequently, driven by this research need, PML has garnered vibrant attention in machine learning [32, 51]. It represents a new paradigm for multi-label recognition (MLR) and has been widely applied in micro-video classification [23], image recognition [8]. In these applications, the training dataset typically contains only partial labels rather than complete ones, which better reflects the complexity of real-world data. Unlike standard partial label learning, which selects the true label from a candidate set, our setting focuses on completing the missing entries in the label matrix based on known positive and negative annotations. For instance, in a scenario where one instance corresponds to a single label, the model only needs to identify the correct label for that specific instance (Fig. 1 (a)). Conversely, in a setting where multiple instances correspond to a single label, the model must extract features from several instances to determine the common label (Fig. 1 (b)).
@@
The key to addressing PML tasks lies in deeply mining discriminative semantic features and constructing a collaborative representation model to fully leverage both label and instance information. A straightforward approach involves leveraging matrix factorization with auxiliary information to accelerate label completion and enhance prediction performance [46]. However, such works suffer from performance limitations due to their neglect of high-order correlations among labels. This inherent constraint has motivated extensive research on capturing label correlations and exploiting data structural information through various mechanisms. For instance, Xie et al. [45] proposed a partial label framework that assigns confidence scores to candidate labels while utilizing data structural information to jointly optimize the classification model and refine confidence estimations for ground-truth labels. Concurrently, semantic-aware modules [3, 27] and loss calibration techniques [50] have been widely adopted in partial multi-label frameworks due to their competitive performance in exploring label correlations.
@@
To date, visual-language pre-trained models have exhibited outstanding performance across various visual tasks. These models, such as the CLIP [29], pre-trained on a vast corpus of image-text pairs, benefit from large-scale pre-training and effectively bridge the gap between vision and language, providing rich prior knowledge for downstream tasks. In the context of PML, these pre-trained models have demonstrated significant advantages. Even when annotations are available for only a subset of categories, they can leverage their generalization capabilities and existing knowledge of visual-language associations to accurately infer unseen categories, thereby enhancing recognition accuracy and robustness. Sun et al. [33] utilized the strong alignment capability of pre-trained visual-language models by introducing a pair of learnable positive and negative prompts, along with a category-specific region feature aggregation method, addressing how to train models and improve recognition accuracy under limited annotations effectively. Ding et al. [7] particularly highlighted the importance of leveraging structured semantic relationships between labels to address the lack of supervision in situations of incomplete labeling. By introducing cross-modal prompters and a semantic association module, their method more effectively extracts and utilizes the rich knowledge embedded in pre-trained visual-language models, such as CLIP, resulting in performance improvements on MLR tasks beyond existing approaches. In addition, Wang et al. [35] proposed a hierarchical semantic prompting network using CLIP for hierarchical prompt learning, aimed at multi-label classification tasks with a single positive label. As a result, we have adopted techniques based on multi-modal large language models to tackle PML tasks, thereby enhancing performance and accuracy.
@@
Although the aforementioned methods have achieved satisfactory results to a certain extent, their effectiveness remains constrained in addressing the challenge of insufficient supervision, particularly regarding inadequate exploitation of supervisory signals among instances. In reality, inter-instance relationships can provide crucial clues for understanding and predicting attributes of unlabeled instances. Taking image data as an example, instances within images typically do not exist in isolation but exhibit intrinsic interconnections. Such relational patterns offer additional contextual information that facilitates more accurate target recognition under incomplete label supervision. Moreover, large-scale cross-modal datasets inherently contain rich relational information among instances. While most current multi-modal representation learning approaches have yet to fully exploit these patterns and interactions in PML tasks, they demonstrate significant potential for enhancing model capabilities in comprehending and modeling instance-level associations within complex scenes.
@@
Based on the aforementioned issues, we propose a novel approach to address these challenges. This method begins by deeply excavating the discriminative semantic features embedded within the image context. Subsequently, it constructs a collaborative representation model of labels and instances in a high-dimensional data space, thereby achieving precise mapping across a cross-modal semantic space. Specifically, to tackle the challenges posed by incomplete labels, we focus on matching multiple semantic co-occurrence relationships between labels and instances. By thoroughly analyzing their interaction patterns, the model is capable not only of identifying existing semantic labels within the data but also of inferring unseen labels based on the associations among known instances. Consequently, this significantly enhances the overall performance of PML tasks in complex scenarios and large-scale cross-modal environments, offering a new perspective on addressing the challenge of partial labeling.
@@ II Related Work
This section provides a concise overview of previous research on multi-label image classification tasks, focusing on MLR with comprehensive annotations, PML, and multi-modal representation learning in visual tasks.
@@ II-B Partial Multi-Label Learning
In PML, the objective is to learn from a label matrix with missing entries, where available annotations include both known positive and known negative labels. This scenario is more challenging than traditional multi-label learning with complete label assignments. Specifically, our goal is to recover the statuses of unobserved labels by leveraging semantic co-occurrence, distinct from the disambiguation task in standard partial label learning. Formally, given a training dataset $\mathcal{D}=\{(\mathbf{X}_{i},Y_{i})\}_{i=1}^{N}$, where $\mathbf{X}_{i}\in\mathbb{R}^{d}$ is a d-dimensional feature vector representing the $i$-th image. The label space $\mathcal{Y}$ for the $i$-th instance is partitioned into three disjoint subsets: the set of verified positive labels $\mathcal{Y}^{+}_{i}$, the set of verified negative labels $\mathcal{Y}^{-}_{i}$, and the set of unobserved (missing) labels $\mathcal{Y}^{U}_{i}$. The goal is to leverage the partial supervision from $\mathcal{Y}^{+}_{i}$ and $\mathcal{Y}^{-}_{i}$ to accurately infer the true status of labels in $\mathcal{Y}^{U}_{i}$. Early studies [25] often treated missing labels as negative labels by default. However, this simplistic approach can introduce significant label bias, as it may erroneously classify a substantial number of true positive labels as negative, thereby adversely affecting the model’s learning performance. To address this issue, several methods have been proposed. For instance, some approaches leverage instance similarity, label co-occurrence information, and low-rank regularization to recover missing labels through matrix completion techniques [41]. Kapoor et al. [17] treated missing labels as latent variables within a probabilistic framework and updated them via Bayesian inference. Additionally, Wu et al. [42] introduced a third state for missing labels and reconstructed the complete label information by enforcing consistency constraints and label assignment smoothness constraints. It is noteworthy that some earlier works [1] still adhered to the strategy of treating unknown labels as negative, while the approach of decomposing the multi-label task into multiple independent binary classification problems overlooked the inherent dependencies among labels.
@@
In light of the aforementioned limitations, subsequent research has increasingly focused on leveraging label correlations to supplement unknown label information, employing methods such as low-rank regularization, label dependency propagation, and hybrid graph modeling [46]. However, these approaches typically require loading the entire training dataset into memory, making them difficult to integrate with fine-tuning strategies for deep neural networks based on mini-batch training. To address this, recent studies have shifted their attention to designing novel loss functions and optimization strategies. For instance, Durand et al. [8] proposed a generalization of the standard binary cross-entropy loss by incorporating label proportion information. Huynh and Elhamifar [16] introduced regularization to the loss function by considering statistical co-occurrence and image-level feature similarity. Pu et al. [27] enhanced samples by generating composite labels through the mixing of category-specific features in the feature space. Meanwhile, Xie et al. [44] utilized structural information from both feature and label spaces to assign confidence scores to potential labels, while Wang et al. [36] guided label disambiguation using interior-point methods and gradient boosting algorithms. He et al. [14] designed a soft labeling threshold operator to narrow the gap between candidate labels and ground truth labels. On the other hand, Guan et al. [12] proposed a semi-supervised PML method that combines partially labeled data with unlabeled data, capturing intrinsic label relationships through low-rank and manifold constraints. To address the issue of significant dependencies among labels, Chen et al. [3] inferred missing labels by leveraging known label information within and across images, leading to the development of the HST method [2], which is based on category-specific feature-prototype similarity and differential threshold learning. Although HST incorporates adaptive threshold learning, it still relies on modeling inter-class similarities, which may vary across datasets, potentially affecting generalization performance. These methods primarily rely on co-occurrence information within datasets, leading to suboptimal inference when labels are sparse or unstable. In contrast to existing approaches, our method leverages pre-trained multi-modal models, enabling effective reasoning in low-label scenarios by incorporating prior knowledge. To tackle the issue of incomplete label annotations, we introduce an intrinsic semantic augmentation strategy to ensure robust performance.
@@ II-C Multi-modal Semantic Fusion
Based on this context, researchers have proposed various innovative methods to explore further and leverage the semantic relationships between vision and language. For instance, Wang et al. [35] developed a novel hierarchical semantic prompting network that dynamically captures the hierarchical semantic relationships between images and text, effectively enhancing the model’s ability to understand complex semantic structures. Concurrently, He et al. [15] introduced an open-vocabulary multi-label classification framework based on multi-modal knowledge transfer, aiming to address the issue of unseen labels in real-world MLR systems. Additionally, Wang et al. [39] designed a multi-modal context prompting learning method, which incorporates learnable context prompts in both image and text modalities and utilizes coupling functions to facilitate interactive connections between modal prompts, thereby strengthening the correspondence between images and labels and improving the accuracy of MLR. Liu et al. [21] proposed a novel multi-label few-shot image classification framework based on paired feature enhancement and flexible prompt learning to explore further and leverage the semantic relationship between vision and language. Rawlekar et al. [30] introduced PositiveCoOp, which learns only positive prompts and replaces negative prompt learning with embedded vectors learned directly in the shared feature space without relying on a text encoder. Liu et al. [22] presented a language-driven cross-modal classifier, which trains a classifier using language data generated by large language models and then transfers it to the visual modality. Fang et al. [11] proposed ProPOT, which generates improved class prototypes by partially aggregating local image features through a partially optimal transport scheme.
Despite these advancements, existing multi-modal learning methods still exhibit limitations. They fail to adequately capture the complex co-occurrence relationships between labels and instances, often overlooking cross-dependencies, which adversely affect alignment performance. Moreover, current models tend to focus on local similarities while lacking the integration of global correlations, thereby constraining the accurate estimation of label confidence. To address these issues, our approach employs global modeling of labels, instances, and their multiple co-occurrence relationships, achieving multi-level semantic fusion and significantly enhancing alignment precision. Furthermore, we construct a cross-modal fusion module to deeply integrate textual and visual data, optimizing label confidence estimation. To address the issue of incomplete label annotations, we propose an intrinsic semantic augmentation strategy, which ensures robust performance.
@@ III Methodology
We propose learning and inferring complex semantic co-occurrence relationships between the visual and textual feature spaces, particularly among labels and labeled instances. This approach enhances the model’s ability to identify rare or ambiguous categories. It boosts its generalization skills, enabling it to navigate the complexity and diversity of the real world more effectively. Fig. 2 outlines our proposed SCINet method. Initially, we process and interact with input data using the bi-dominant prompter module. Simultaneously, Learnable Prompts acquire textual labels in the form of a vector sequence, providing context for the given label names. Subsequently, text-dominant and image-dominant encoders process textual and image features, respectively. The cross-modality fusion module deeply combines text and visual information to understand and classify data more effectively. Furthermore, we utilize an intrinsic semantic augmentation strategy to enable the model to comprehend data features and encourage a synergistic relationship between label confidence and sample complexity.
@@ III-A Bi-Dominant Prompter
Advancements in large-scale pre-trained models, such as BERT and CLIP, facilitate efficient access to contextual label embeddings. This advancement enables the effortless inference of co-occurrence relationships or correspondences between labels and instances. Such a strategy is especially attractive when adequate label supervision is lacking. Furthermore, the extensive prior knowledge contained within pre-trained models facilitates the connection between known and unknown labels, potentially alleviating the challenges associated with limited label supervision. Therefore, we introduce a semantics-driven prompter to explore the label-to-label co-occurrence relationships inherent in the pre-trained model. Given the widespread attention and outstanding performance of vision-language models in computer vision, we have chosen CLIP as the primary focus of our research. Let $V=[v_{1},v_{2},\cdots,v_{m},\text{CLS}]$, where each $v_{i}$ represents a learnable soft prompt token, and CLS denotes the given label name. The entire set of labels is denoted as $Y$, representing each instance as $y_{i}\in Y$. The text-dominant encoder accepts label information while simultaneously processing the image set. The textual representation obtained from the text-dominant encoder process is denoted as $z$. The input image $\mathbf{X}$ is represented as $\mathbf{X}=\left\{\left(\mathbf{s_{1}},C_{1},U_{1}\right),\left(\mathbf{s_{2}},C_{2},U_{2}\right),\cdots,\left(\mathbf{s_{n}},C_{n},U_{n}\right)\right\}$, where $s_{i}$ denotes the $i$-th instance, $C_{i}=\{\mathbf{c}_{i1},\mathbf{c}_{i2},\cdots\}$ represents its set of known correct labels, and $U_{i}=\{\mathbf{u}_{i1},\mathbf{u}_{i2},\cdots\}$ corresponds to its set of unknown labels. To enhance the model’s robustness, three different levels of transformations are applied to the input, including weak transformation ($X^{-}$), original transformation ($X$), and strong transformation ($X^{+}$). These three transformations are collectively integrated as $X^{{}^{\prime}}$ and utilized for subsequent learning and optimization processes. For a detailed explanation of the three different levels of transformation, please refer to Section III-C. The visual representation obtained, denoted as $f$, is extracted through the modified CLIP image encoder. This dual-encoder approach enables the model to understand and apply the semantic co-occurrence knowledge, thereby improving the performance of MLR. Specifically, the text-dominant encoder and the visual-dominant encoder are defined as follows: